OpenStudy (anonymous):

(10^25)-7 is divisible by a) 2 b)3 c)9 d) both 2 & 3

OpenStudy (anonymous):

b) 3

OpenStudy (anonymous):

this will be and odd so wont be divisible by 2 it will consist of 24 9's and end in 3 so it is divisible by 3

OpenStudy (anonymous):

please explain i can't understand

OpenStudy (anonymous):

10^2 - 7 = 100 - 7 = 93 one 1 nine followed by 3 10^3 - 7 = 1000 - 7 = 993 2 nines followed by 3 10^4 - 7 = 10000 - 7 = 9993 3 nines followed by 3 following the above pattern we find that 10^25 - 7 will have 24 nines with 3 at the end so it will be an odd no so 2 is ruled out so a) 2 and d) both 2 & 3 are ruled out for 3 and 9 we have to find sum of all the digits it will b (9x24) + 3 = 219 and 219 is only divisible by 3 and not by 9 so correct ans is b)3 hope it is clear now ☺

OpenStudy (anonymous):

10^25-7=(10^25-1)-6 the no 10^25-1 =999.....99 (25 digits) is divisible by 3 & 9 therefore (10^25-1)-6=(24 nines & unit digit is 3) i.e. 999......93 this number is only divisible by 3.